OpenStudy (kpop4life123):

The equation shows the relationship between x and y: y = –8x + 3 What is the slope of the equation? (5 points) –8 –5 3 8

RhondaSommer (rhondasommer):

y=mx + b

RhondaSommer (rhondasommer):

m= slope y= -8x + 3 y= mx + b

RhondaSommer (rhondasommer):

what is slope?

OpenStudy (kpop4life123):

i dont actually know :/

RhondaSommer (rhondasommer):